OpenStudy (anonymous):

Find functions f and g so that h(x) = (f ∘ g)(x). h(x) = (6x - 14)8 A. f(x) = 6x - 14, g(x) = x8 B. f(x) = 6x8 - 14, g(x) = -14 C. f(x) = x8, g(x) = 6x - 14 D. f(x) = (6x)8, g(x) = -14

OpenStudy (jdoe0001):

so got any suspects in mind?

OpenStudy (jdoe0001):

care to try any of the choices? which one do you think we could try?

OpenStudy (anonymous):

Tbh I have no idea how to do this, but I would love to learn how to!

OpenStudy (jdoe0001):

ok... gimme a choice to try to see if it works

OpenStudy (anonymous):

B. f(x) = 6x8 - 14, g(x) = -14

OpenStudy (jdoe0001):

B, ok let's see if B works \(\bf h(x) = (6x - 14)^8 \\ \quad \\ \quad \\ (B) \qquad f(x) = 6x^8 - 14\qquad {\color{red}{ g(x)}} = -14 \\ \quad \\ (f \ o \ g)(x)\implies f(\quad g(x)\quad )=6({\color{red}{ g(x)}})^8 - 14\implies 6({\color{red}{ -14}})^8 \\ \quad \\ \quad \\ 6({\color{red}{ -14}})^8 - 14 \ne (6x - 14)^8\) so no dice on that one what about another choice?

OpenStudy (anonymous):

D. f(x) = (6x)8, g(x) = -14

OpenStudy (jdoe0001):

\(\bf h(x) = (6x - 14)^8 \\ \quad \\ \quad \\ (D) \qquad f(x) = (6x)^8 - 14\qquad {\color{red}{ g(x)}} = -14 \\ \quad \\ (f \ o \ g)(x)\implies f(\quad g(x)\quad )=[6({\color{red}{ -14}})]^8 - 14 \\ \quad \\ [6({\color{red}{ -14}})]^8 - 14 \ne (6x - 14)^8\) no dice on that one either what about another one?

OpenStudy (jdoe0001):

hmm actaully it has no 14.. lemme fix that

OpenStudy (jdoe0001):

\((\bf h(x) = (6x - 14)^8 \\ \quad \\ \quad \\ (D) \qquad f(x) = (6x)^8 \qquad {\color{red}{ g(x)}} = -14 \\ \quad \\ (f \ o \ g)(x)\implies f(\quad g(x)\quad )=[6({\color{red}{ -14}})]^8 \\ \quad \\ [6({\color{red}{ -14}})]^8 \ne (6x - 14)^8\) so... no dice anyhow another choice?

OpenStudy (anonymous):

Is choice (B) transcribed correctly?

OpenStudy (jdoe0001):

yes, keep in mind that g(x) is just a constant

OpenStudy (anonymous):

C. f(x) = x8, g(x) = 6x - 14

OpenStudy (anonymous):

no, choice (B)

OpenStudy (jdoe0001):

C, let's try that one then \(\bf f(x) = x8, g(x) = 6x - 14 \\ h(x) = (6x - 14)^8 \\ \quad \\ \quad \\ (D) \qquad f(x) = x^8 \qquad {\color{red}{ g(x)}} = 6x -14 \\ \quad \\ (f \ o \ g)(x)\implies f(\quad g(x)\quad )=({\color{red}{ 6x -14}})^8 \\ \quad \\ ({\color{red}{ 6x -14}})^8 \iff (6x - 14)^8\)

OpenStudy (jdoe0001):

got the wrong letter choice anyhow \(\bf h(x) = (6x - 14)^8 \\ \quad \\ \quad \\ (C) \qquad f(x) = x^8 \qquad {\color{red}{ g(x)}} = 6x -14 \\ \quad \\ (f \ o \ g)(x)\implies f(\quad g(x)\quad )=({\color{red}{ 6x -14}})^8 \\ \quad \\ ({\color{red}{ 6x -14}})^8 \iff (6x - 14)^8\)

OpenStudy (anonymous):

So C is correct?

OpenStudy (jdoe0001):

f ( g(x) ) <---- as you can see, means, g(x) will replace any "x" in the function f(x) so yes, it's C

OpenStudy (anonymous):

Gotcha! (: Thank you very much

OpenStudy (jdoe0001):

yw